### v2

@@ 4.2 Verification and Falsification Feedback
Leveraging the analytical nature of candidate Lyapunov functions, efficient methods for symbolic expressions, such as root finding Feng et al. (2024), can be applied to Lyapunov condition verification and counterexample generation. In this work, we propose a global-optimization-based numerical verification algorithm, using Simplicial Homology Global Optimization (SHGO) (Endres et al., 2018), that effectively checks the Lyapunov conditions around minimizers and feedback counterexamples into the training set $\mathcal{X}$ for reward calculation. SHGO is a constrained global optimization algorithm with theoretical guarantees for convergence. To make the paper self-contained, we present the guarantees in Proposition 4.1. This algorithm identifies the global minimizer over the state space from a set of local minimums, each of which is obtained from a convex sub-domain in the feasible search space. Taking advantage of the theoretical results, we employ the SHGO algorithm for counterexample detection in our verification process.
@@ 4.4 Automated Expert Guidance
The GP algorithm starts with a batch of initial populations (expression trees) and iteratively refines these populations through evolutionary operations: mutation, selection, and crossover, with a pre-defined metric to evaluate the fitness of populations to the task (e.g. MSE for symbolic regression task). Within our framework, we feed the latest batch of generated candidates $\{\tilde{V}_{\phi}^{i}\sim p(\tilde{V}_{\phi}|\phi,f(x))\}_{i=1}^{Q}$ from the decoder into the GP module as the starting population [^1] Without a good initial population from the transformer, GP algorithms itself face significant challenges in directly finding valid Lyapunov functions for high-dimensional systems due to the exponentially growing search space. Appendix H.3., refine these expressions through evolutionary operations with Lyapunov risk reward as the measure of fitness, and obtain a batch of refined expressions. We select an ‘elite set’ of the refined expressions $\tilde{\mathcal{V}}_{gp}=\{\tilde{V}_{gp}^{i}\sim\text{GP}(\tilde{V}_{\phi})\}_{i=1}^{G}$, regard them as target expressions, and optimize the transformer model in a supervised learning manner, maximizing the probability that the generated token matches the reference tokens from $\tilde{V}_{gp}$, with the following expert guidance loss:
@@
where $G$ is the number of expressions in $\tilde{\mathcal{V}}_{gp}$, and $k_{i}$ is the complexity (number of symbolic tokens in the pre-order traversal) of $\tilde{V}_{gp}^{i}$. Each expression $\tilde{V}_{gp}^{i}$ is weighted by its Lyapunov risk reward in $\mathcal{L}$. Algorithm 1 summarizes the training process, with more details in Appendix D. The GP solutions explore the characteristics of Lyapunov functions that have not been captured by the transformer yet and effectively guide the transformer for higher-quality generation.
@@ 5 Experiment
We validate the proposed algorithm across a variety of nonlinear dynamics by finding their local Lyapunov functions at the equilibrium point to verify their stability, where the systems are autonomous (or closed-loop systems with known feedback control laws). We use dReal Gao et al. (2013) SMT solver for final verification of found Lyapunov functions, with a numerical tolerance error $\epsilon=e^{-3}$ and precision $\delta=e^{-12}$, over the state space, i.e. $V(x)>\delta$ and $L_{f}V(x)<-\delta$ over $\mathcal{D}\backslash\mathcal{B}_{\epsilon}(0)$. The excluded ball $\mathcal{B}_{\epsilon}(0)$ is to avoid numerical issues, which is a common practice for SMT-based formal verification Chang et al. (2019). Global stability can be determined through further analysis; for instance, if the Lie derivative is a negation of sum-of-squares, it is sufficient to establish global stability.
@@ 5.1 Experimental Setting
Test Dynamics. We categorize our collected test dynamical systems into two kinds: 1). Polynomial Systems, and 2). Non-polynomial Systems, where three polynomial systems are adopted from Alfarano et al. (2024) (Appendix F.2, F.3) and others come from real-world examples. Detailed information about systems is summarized in Appendix F and G. The dimension of these test systems ranges up to 10.
Implementation Details of Framework. Detailed explanation for all components in our framework is presented in Appendix A (Transformer), B (Global-optimization-based Numerical Verification), C (Risk-seeking Policy Gradient), and D (Genetic Programming). The symbolic library $\mathcal{L}_{s}$ is defined as {+,,×$\{+,-,\times,\sin,\cos,x_{i}\}$ in all tests.
Baseline Algorithms. We compare our proposed framework against two learning-based algorithms for neural Lyapunov function discovery in continuous nonlinear dynamical systems: 1) Augmented Neural Lyapunov Control (ANLC) Grande et al. (2023), and 2) FOSSIL 2.0 Edwards et al. (2024). Both works train a neural Lyapunov function with the empirical Lyapunov risk loss and employ a counter-example guided inductive synthesis (CEGIS) loop for better generalization over the state space $\mathcal{D}$. Detailed descriptions and baseline setup can be found in Appendix E. We also directly apply the pre-trained model in Alfarano et al. (2024) to our test systems. Limited to global stability and systems with dimension less than $6$, their model can only succeed in systems presented in Appendix F.1, F.2, F.3, and G.1.
Table 2: Training time and success rate comparison between proposed framework and baseline algorithms. The Succ % is the successful rate of finding a valid Lyapunov function within 5 hours out of 5 random seeds. Runtime is the average training time for a successful trial.
@@ 5.2 Performance Analysis
Analytical Lyapunov functions can potentially bypass the need for formal verification. In the 3-D Trig system (Appendix G.2), over the state space $\mathcal{D}=\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}\bm{\mid}|x_{i}|\leq 1.5,\forall\>i\in\{1,2,3\}\}$, the positive definiteness of the identified Lyapunov function is evident from its formulation. Moreover, the Lie derivative $L_{f}V=-2x_{2}^{2}-x_{3}\sin(2x_{3})$ is directly identifiable as non-positive in $\mathcal{D}$, since $x\sin(x)>0$ for all $x\in(-\pi,\pi)$. By the invariance principle Khalil (2002), the discovered function certifies the asymptotic stability of origin in state space $\mathcal{D}$. When direct identification is non-trivial, SMT solvers can efficiently verify Lyapunov conditions given analytical formulations’ simplicity.
@@ 5.4 Comparisons with Baselines
Table 2 compares our success rate and training runtime with two baselines across various test dynamics. For low-dimensional systems, both baselines achieve notably shorter overall training runtime. However, to achieve efficient verification, these methods rely on relatively small neural networks, which can fail to converge on more challenging tasks (e.g., the simple pendulum and 3-D Trig dynamics involving trigonometric terms). Consequently, both baselines exhibit lower overall success rates than our approach.
@@ 5.5 Ablation Studies
Verification Comparison. We compare three verification methods—(1) root-finding (Feng et al., 2024), (2) the dreal SMT solver, and (3) random sampling—against our approach on the 6-D polynomial system. SHGO-based counterexample feedback is an adversarial reward, allowing fast refinement and stronger final guarantees but risking training instability. In contrast, random sampling yields smoother rewards and more stable training, yet it identifies violations less efficiently. We thus blend both methods to balance final performance and training stability. Meanwhile, the SMT solver and root-finding produce mostly mild violations, offering limited optimization guidance (Appendix H.2).
@@ 6 Conclusion
This paper presents an end-to-end framework for analytical Lyapunov function discovery on nonlinear dynamical systems, which employ a symbolic transformer for candidate expressions’ generation, with risk-seeking policy gradient together with GP for transformer training. We propose an efficient verification and counter-example generation method in the training process via a global optimization algorithm SHGO, and use SMT solvers for the final Lyapunov function certification. Our framework can be extended to other certificate-function discoveries, such as deriving control barrier functions for safety in given dynamics. More broadly, several promising research directions arise from this work. One direction is efficiently incorporating physical constants such as gravitational acceleration or object mass, which can significantly improve generalization—though directly introducing them as variables might add unnecessary complexity. Another direction involves extending our framework to discover control Lyapunov functions, which are vital for designing stabilizing feedback laws. Finally, since constructing large-scale datasets for local Lyapunov functions remains challenging, leveraging our approach to refine pre-trained models for downstream tasks represents an exciting opportunity to enhance Lyapunov function discovery.
Impact Statements
@@ Appendix A Symbolic Transformer Model
We outline the details of our symbolic transformer model, a conditional generator for analytical candidate expression generation, comprising two components: 1) encoder, and 2) decoder. $\phi=\{\zeta,\theta\}$ denotes the transformer parameters, where $\zeta$ and $\theta$ denote the parameters of encoder and decoder respectively.
@@ Appendix B Global-optimization-based Numerical Verification
For a given dynamics $f(x)$, suppose $\tilde{V}_{\phi}$ is an invalid analytical candidate Lyapunov function. According to Lyapunov conditions defined in Proposition 3.4, for $x_{1}^{*},x_{2}^{*}\in\mathcal{D}$, where $x_{1}^{*},x_{2}^{*}$ are the global minimizers of $\tilde{V}_{\phi}$ and $-L_{f}\tilde{V}_{\phi}$ in the state space $\mathcal{D}$, the following two inequalities hold: $\tilde{V}_{\phi}(x_{1}^{*})\leq 0$ and $L_{f}\tilde{V}_{\phi}(x_{2}^{*})\geq 0$. This implies that if $\tilde{V}_{\phi}$ is invalid, the neighborhoods of $x_{1}^{*}$ and $x_{2}^{*}$ are highly likely to capture significant violations. Based on this observation, we propose a global-optimization-based numerical verification. This verification identifies minimizers $x_{1}^{*}$ and $x_{2}^{*}$ by Simplicial Homology Global Optimization (SHGO), verifies Lyapunov conditions on localized samples in neighborhoods $\mathcal{B}_{r}(x_{1}^{*})$ and $\mathcal{B}_{r}(x_{2}^{*})$, and feeds counterexamples back into the training set $\mathcal{X}$. We detail the sampling and condition-checking procedures in Algorithm 2. Figure 4 illustrates this verification process on a sampled candidate, $\tilde{V}_{\phi}=(x_{1}+x_{2})^{2}+x_{2}$, for the Van der Pol Oscillator. In the implementation, we initiate with $2048$ starting points and iterate $3$ times in the SHGO algorithm for the minimizer detection. We the tested number of starting points with values $[1024,2048,4096,8196]$ on various high dimensional continuous functions, and the setting with $2048$ starting points achieves the best efficiency. In datapoint sampling for counter-example identification, for each candidate expression, $800$ data points are sampled from each of $\mathcal{B}_{r}(x_{1}^{*})$ and $\mathcal{B}_{r}(x_{2}^{*})$, and additional $800$ data points are randomly sampled across the state space $\mathcal{D}$.
Figure 4: This plot visualizes our proposed verification process on a sampled candidate. $\tilde{V}_{\phi}=(x_{1}+x_{2})^{2}+x_{2}$ is a sampled candidate during the training for Van Der Pol Oscillator. Using Simplicial Homology Global Optimization, we first identify the minimizer of the Lyapunov function and the maximizer of the Lie Derivative, the black dots in each graphs. Next, data points are sampled in the neighborhoods of the two points, the regions in red circles. For sampled data points violate the Lyapunov conditions, we feed it into the training set $\mathcal{X}$.

### v3

@@ 4.2 Verification and Falsification Feedback
Leveraging the analytical nature of candidate Lyapunov functions, efficient methods for symbolic expressions, such as root finding Feng et al. (2024c), can be applied to Lyapunov condition verification and counterexample generation. In this work, we propose a global-optimization-based numerical verification algorithm, using Simplicial Homology Global Optimization (SHGO) (Endres et al., 2018), that effectively checks the Lyapunov conditions around minimizers and feedback counterexamples into the training set $\mathcal{X}$ for reward calculation. SHGO is a constrained global optimization algorithm with theoretical guarantees for convergence. To make the paper self-contained, we present the guarantees in Proposition 4.1. This algorithm identifies the global minimizer over the state space from a set of local minimums, each of which is obtained from a convex sub-domain in the feasible search space. Taking advantage of the theoretical results, we employ the SHGO algorithm for counterexample detection in our verification process.
@@ 4.4 Automated Expert Guidance
The GP algorithm starts with a batch of initial populations (expression trees) and iteratively refines these populations through evolutionary operations: mutation, selection, and crossover, with a pre-defined metric to evaluate the fitness of populations to the task (e.g., MSE for symbolic regression task). Within our framework, we feed the latest batch of generated candidates $\{\tilde{V}_{\phi}^{i}\sim p(\tilde{V}_{\phi}|\phi,f(x))\}_{i=1}^{Q}$ from the decoder into the GP module as the starting population [^1] Without a good initial population from the transformer, GP algorithms itself face significant challenges in directly finding valid Lyapunov functions for high-dimensional systems due to the exponentially growing search space. Appendix H.3., refine these expressions through evolutionary operations with Lyapunov risk reward as the measure of fitness, and obtain a batch of refined expressions. We select an ‘elite set’ of the refined expressions $\tilde{\mathcal{V}}_{gp}=\{\tilde{V}_{gp}^{i}\sim\text{GP}(\tilde{V}_{\phi})\}_{i=1}^{G}$, regard them as target expressions, and optimize the transformer model in a supervised learning manner, maximizing the probability that the generated token matches the reference tokens from $\tilde{V}_{gp}$, with the following expert guidance loss:
@@
where $G$ is the number of expressions in $\tilde{\mathcal{V}}_{gp}$, and $k_{i}$ is the complexity (number of symbolic tokens in the pre-order traversal) of $\tilde{V}_{gp}^{i}$. Each expression $\tilde{V}_{gp}^{i}$ is weighted by its Lyapunov risk reward in $\mathcal{L}$. Algorithm 1 summarizes the training process, with more details in Appendix D. The GP solutions explore the characteristics of Lyapunov functions that have not been captured by the transformer yet and effectively guide the transformer.
Remark 4.2 (Exploration–exploitation trade-off).
In our framework, exploration arises during candidate expression generation and through GP’s evolutionary operations, while exploitation is driven by the risk-seeking policy gradient and the expert-guided loss.
@@ 5 Experiment
We validate the proposed algorithm across a variety of nonlinear dynamics by finding their local Lyapunov functions at the equilibrium point to verify their stability, where the systems are autonomous (or closed-loop systems with known feedback control laws). We use dReal Gao et al. (2013) SMT solver for final verification of found Lyapunov functions, with a numerical tolerance error $\epsilon=e^{-3}$ and precision $\delta=e^{-12}$, over the state space, i.e. $V(x)>\delta$ and $L_{f}V(x)<-\delta$ over $\mathcal{D}\backslash\mathcal{B}_{\epsilon}(0)$. The excluded ball $\mathcal{B}_{\epsilon}(0)$ is to avoid numerical issues, which is a common practice for SMT-based formal verification (Chang et al., 2019). Global stability can be determined through further expert analysis; for instance, if the Lie derivative is a negation of SOS, it is sufficient to establish global stability.
@@ 5.1 Experimental Setting
9-D Synthetic Sys (App. G.6) 27047s $6.6$ s $\begin{array}[t]{l}V=\left(\sum_{i=1}^{6}x_{i}^{2}\right)+\sin(x_{7})^{2}\\ +x_{8}^{2}-\cos(x_{9})+1\end{array}$ l.a.s. 60
@@
Test Dynamics. We categorize our collected test dynamical systems into two kinds: 1). Polynomial Systems, and 2). Non-polynomial Systems, where three polynomial systems are adopted from Alfarano et al. (2024) (Appendices F.2 & F.3), and others come from real-world examples. Detailed information about systems is summarized in Appendices F and G. The dimension of these test systems ranges up to 10.
Implementation Details of Framework. Detailed explanation for all components in our framework is presented in Appendices A (Transformer), B (Global-optimization-based Numerical Verification), C (Risk-seeking Policy Gradient), and D (Genetic Programming). The symbolic library $\mathcal{L}_{s}$ is defined as {+,,×$\{+,-,\times,\sin,\cos,x_{i}\}$ in all tests.
Baseline Algorithms. We compare our proposed framework against four baseline algorithms for continuous nonlinear dynamics. Neural methods: 1) Augmented Neural Lyapunov Control (ANLC) Grande et al. (2023), and 2) FOSSIL 2.0 Edwards et al. (2024). Analytical methods: 3) the transformer-based global Lyapunov search of Alfarano et al. (2024), and 4) SOS methods via SOSTOOLS (Matlab) (Papachristodoulou et al., 2013). We use the formulation of Papachristodoulou & Prajna (2005a) on polynomial systems, and apply the recasting technique of Papachristodoulou & Prajna (2005b) to convert non-polynomial dynamics into rational form so they can also be handled by SOS.
Both ANLC and FOSSIL 2.0 train a neural Lyapunov function with the empirical Lyapunov risk loss and employ a counter-example guided inductive synthesis (CEGIS) loop for better generalization over the state space $\mathcal{D}$. Alfarano et al. (2024) pre-trains a transformer on global Lyapunov function datasets with beam search for candidate generation. SOS methods formulate Lyapunov functions as the feasible solutions of some semi-definite programming tasks and solve these tasks via convex optimization tools. Details of each baseline can be found in Appendix E.
Table 2: Training time and success rate comparison between ours and the neural baselines. The Succ % is the successful rate of finding a valid Lyapunov function within 5 hours out of 5 random seeds. Runtime is the average training time for a successful trial.
@@ 5.2 Performance Analysis
Analytical Lyapunov functions can potentially bypass the need for formal verification. In the 3-D Trig system (Appendix G.2), over the state space $\mathcal{D}=\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}\bm{\mid}|x_{i}|\leq 1.5,\forall\>i\in\{1,2,3\}\}$, the positive definiteness of the identified Lyapunov function is evident from its formulation. Moreover, the Lie derivative $L_{f}V=-2x_{2}^{2}-x_{3}\sin(2x_{3})$ is directly identifiable as non-positive in $\mathcal{D}$, since $x\sin(x)>0$ for all $x\in(-\pi,\pi)$. By the invariance principle Khalil (2002), the discovered function certifies the asymptotic stability of the origin in state space $\mathcal{D}$. When direct identification is non-trivial, SMT solvers can efficiently verify Lyapunov conditions given analytical formulations’ simplicity.
@@ 5.4 Comparisons with Baselines
Table 3: Training/solving time of ours and sum-of-squares (SOS) on polynomial systems, averaged over successful trials. The stable regions (local or global) of the considered dynamics are indicated in smaller font.
Test Systems 2-D Systems (App. F.1, local; App. F.2, global.) 3-D Systems (App. F.3, global.) 6-D System (App. F.4, local.) 8-D System (App. F.5, local.) 10-D System (App. F.6, local.)
Ours 97s 108s 1667s 14358s 64223s
Sum-of-squares 0.765s 1.503s - - -
Neural Lyapunov Function Baselines. Table 2 compares our success rate and training runtime with two neural Lyapunov function baselines across various test dynamics. For low-dimensional systems, both baselines achieve notably shorter overall training runtime. However, to achieve efficient verification, these methods rely on relatively small neural networks, which fail to converge on more challenging tasks (e.g., the simple pendulum and 3-D Trig dynamics involving trigonometric terms). Consequently, both baselines exhibit lower overall success rates than our approach.
@@
Analytical Lyapunov Function Baselines. We evaluated the pre-trained model of Alfarano et al. (2024) on our benchmarks. Trained solely on globally stable systems with fewer than six states, it produced valid Lyapunov functions only for the low-dimensional examples in Appendices F.1, F.2, F.3, & G.1, which have global stability guarantees, and failed on every benchmark that is only locally stable.
Table 3 compares the training and solving times of our method and the SOS approach on polynomial systems from Appendix F. For low-dimensional systems, SOS finds valid Lyapunov functions more efficiently, with shorter runtimes. However, it fails to certify stability for the higher-dimensional systems in Appendices F.4, F.5, and F.6, which are only locally stable. While SOS methods can scale to 10-dimensional or higher systems for global stability verification, local stability requires additional constraints on the Lie derivative, which significantly impact scalability. For example, in the 6D polynomial system from Appendix F.4, verifying local stability with a degree-2 polynomial candidate introduces hundreds of symbolic terms in the Lie derivative constraint, rendering the problem intractable for SOSTOOLS. More details are provided in Appendix E.4.
To use SOS on non-polynomial systems without relaxing the dynamics, we follow the recasting scheme of Papachristodoulou & Prajna (2005b), replacing each non-polynomial term with auxiliary variables linked by extra (in)equalities. We applied this procedure to the simple pendulum and the 3-D trig system (Appendices G.1 & G.2). SOS did recover Lyapunov certificates in both cases, but at a much higher computational cost than our framework for the 3-D trig example. Three main limitations are revealed during implementation: (i) recasting demands substantial domain expertise and hand-crafted constraints; (ii) added variables and relations greatly increase the computational burden; and (iii) the formulation certifies only stability, not asymptotic stability. For the locally stable 3-D trig system, eliminating the trigonometric terms introduces four auxiliary variables and two equality constraints, pushing the solve time beyond one hour, versus 157s for our method. These limitations undermine SOS’s practicality for high-dimensional, non-polynomial dynamics.
Table 4: Training/solving time of ours and sum-of-squares (SOS) on two non-polynomial systems, averaged over successful trials.
Frameworks App. G.1 App. G.2
Ours 288s 157s
Sum-of-squares 19.58s 6163s
@@ 5.5 Ablation Studies
Verification Comparison. We compare three verification methods—1) root-finding (Feng et al., 2024c), 2) the dreal SMT solver, and 3) random sampling—against our approach on the 6-D polynomial system. SHGO-based counterexample feedback is an adversarial reward, allowing fast refinement and stronger final guarantees but risking training instability. In contrast, random sampling yields smoother rewards and more stable training, yet it identifies violations less efficiently. We thus blend both methods to balance final performance and training stability. Meanwhile, the SMT solver and root-finding produce mostly mild violations, offering limited optimization guidance (Appendix H.2).
@@ 6 Conclusion
This work introduces an end-to-end framework for discovering analytical Lyapunov functions for nonlinear dynamical systems. A symbolic transformer, trained with a risk-seeking policy gradient and augmented by genetic programming, proposes candidate expressions; the SHGO global optimizer rapidly verifies them and generates counterexamples during training; and an SMT solver certifies the final Lyapunov candidates. The framework scales to 10-dimensional dynamics and has discovered previously unknown local Lyapunov functions for lossy power system dynamics. It can be extended to other certificate-function discoveries, such as control barrier functions for safety certificates.
Several promising future research directions emerge from this work. One is efficiently incorporating physical constants, such as the gravitational constant and object mass, which could significantly improve generalization—though directly introducing them as variables may add unnecessary complexity. Another is extending the framework to discover control Lyapunov functions, which are essential for designing stabilizing feedback laws. Theoretical analysis of the framework, including proofs of completeness and convergence, would also be valuable. Finally, since constructing large-scale datasets for local Lyapunov functions remains challenging, using our approach to refine pre-trained models Alfarano et al. (2024) offers an exciting opportunity to further advance Lyapunov function discovery.
Acknowledgment
The authors thank Dr. Huan Zhang at the University of Illinois Urbana-Champaign for helpful discussions. J. Feng is supported by the UC–National Laboratory In-Residence Graduate Fellowship L24GF7923. Y. Shi acknowledges support from NSF ECCS-2200692, DOE grant DE-SC0025495, and the Schmidt Sciences AI2050 Early Career Fellowship.
Impact Statement
@@ Appendix A Symbolic Transformer Model
We outline the details of our symbolic transformer model, a conditional generator for analytical candidate expression generation, comprising two components: 1) an encoder, and 2) a decoder. $\phi=\{\zeta,\theta\}$ denotes the transformer parameters, where $\zeta$ and $\theta$ denote the parameters of encoder and decoder respectively.
@@ Appendix B Global-optimization-based Numerical Verification
For a given dynamics $f(x)$, suppose $\tilde{V}_{\phi}$ is an invalid analytical candidate Lyapunov function. According to Lyapunov conditions defined in Proposition 3.4, for $x_{1}^{*},x_{2}^{*}\in\mathcal{D}$, where $x_{1}^{*},x_{2}^{*}$ are the global minimizers of $\tilde{V}_{\phi}$ and $-L_{f}\tilde{V}_{\phi}$ in the state space $\mathcal{D}$, the following two inequalities hold: $\tilde{V}_{\phi}(x_{1}^{*})\leq 0$ and $L_{f}\tilde{V}_{\phi}(x_{2}^{*})\geq 0$. This implies that if $\tilde{V}_{\phi}$ is invalid, the neighborhoods of $x_{1}^{*}$ and $x_{2}^{*}$ are highly likely to capture significant violations. Based on this observation, we propose a global-optimization-based numerical verification. This verification identifies minimizers $x_{1}^{*}$ and $x_{2}^{*}$ by Simplicial Homology Global Optimization (SHGO), verifies Lyapunov conditions on localized samples in neighborhoods $\mathcal{B}_{r}(x_{1}^{*})$ and $\mathcal{B}_{r}(x_{2}^{*})$, and feeds counterexamples back into the training set $\mathcal{X}$. We detail the sampling and condition-checking procedures in Algorithm 2. Figure 4 illustrates this verification process on a sampled candidate, $\tilde{V}_{\phi}=(x_{1}+x_{2})^{2}+x_{2}$, for the Van der Pol Oscillator. In the implementation, we initiate with $2048$ starting points and iterate $3$ times in the SHGO algorithm for the minimizer detection. We tested the number of starting points with values $[1024,2048,4096,8196]$ on various high-dimensional continuous functions, and the setting with $2048$ starting points achieves the best efficiency. In datapoint sampling for counter-example identification, for each candidate expression, $800$ data points are sampled from each of $\mathcal{B}_{r}(x_{1}^{*})$ and $\mathcal{B}_{r}(x_{2}^{*})$, and additional $800$ data points are randomly sampled across the state space $\mathcal{D}$.
Figure 4: This plot visualizes our proposed verification process on a sampled candidate. $\tilde{V}_{\phi}=(x_{1}+x_{2})^{2}+x_{2}$ is a sampled candidate during the training for Van Der Pol Oscillator. Using Simplicial Homology Global Optimization, we first identify the minimizer of the Lyapunov function and the maximizer of the Lie Derivative, the black dots in each graph. Next, data points are sampled in the neighborhoods of the two points, the regions in red circles. For sampled data points that violate the Lyapunov conditions, we feed them into the training set $\mathcal{X}$.